# Designing lattice spin models and magnon gaps with supercurrents

### v1

@@ Abstract
The fact that the intrinsic magnon gap in the antiferromagnetic insulator is enhanced by supercurrent for one magnon species but suppressed for the other lifts their degeneracy. Interestingly, this is the key to inducing magnon spin currents in antiferromagnets via for instance the spin Seebeck effect Rezende et al. (2019).
Our results thus demonstrate a route to electrically tunable magnon spin currents in antiferromagnetic insulators, controlled by a dissipationless supercurrent. With regard to the magnitude of the supercurrent-induced magnon gap, our approach is perturbative in nature and thus only small modifications of the magnon gap are within the regime of validity of our methodology. However, a non-perturbative approach using stronger coupling would very likely increase this magnitude, and thus experiments probing this effect in bilayers of superconductors and antiferromagnets with good interfacial coupling would be of high interest. Finally, we underline that our results apply to an altermagnetic insulator as well by generalizing the model to 2D.
Summarizing, we predict that charge and spin supercurrents in superconductors enable electric control over both magnon gaps and magnetic interactions in a way that permits controllable designing of a spin lattice. This reveals a synergy between superconductivity and magnetism that opens up a research avenue with relevance to applications where spin interactions are the key ingredient.
@@ Acknowledgements.
This work was supported by the Research Council of Norway through Grant No. 323766 and its Centres of Excellence funding scheme Grant No. 262633 “QuSpin.” The numerical calculations were performed on resources provided by Sigma2—the National Infrastructure for High Performance Computing and Data Storage in Norway, project NN9577K.
@@ I.1 Diagonalizing the Hamiltonian
The mean-field Hamiltonian for a triplet superconductor with equal-spin Cooper pairs is
@@ I.2 Effective spin-spin interaction
The coupling between the itinerant electrons and the classical spins is given by
@@
The goal is to perform a Schrieffer-Wolff transformation to get an effective spin model. The full Hamiltonian is $H=H_{c}+H_{\text{SC}}$, where we treat $H_{c}$ as a perturbation. Note that diagonal elements of the perturbation $H_{c}$ should be absorbed into $H_{\text{SC}}$ in order to allow for first-order contributions in the transformed Hamiltonian. In other words, if $\langle H_{c}\rangle_{\gamma}\neq 0$, these terms should be included in $H_{\text{SC}}$. $\langle H_{c}\rangle_{\gamma}$ is calculated in Eq. (S17), and in the presence of a pure charge or spin supercurrent, this term is zero. For a spin-polarized charge supercurrent, the first-order term, which is effectively a magnetic field, will dominate the spin Hamiltonian. From here on, we assume that the interaction between the magnetic adatoms is induced by a pure charge or spin supercurrent, and we calculate the spin-spin interaction to second order in $\Lambda$. The Hamiltonian is canonically transformed to $H^{\prime}=\text{e}^{\text{i}S}H\text{e}^{-\text{i}S}$, where $S$ is a hermitian operator known as a generator. This transformed Hamiltonian has the same spectrum as the original Hamiltonian. If the generator is chosen such that $[S,H_{\text{SC}}]=\text{i}H_{c}$, the transformed Hamiltonian is $H^{\prime}=H_{\text{SC}}+\frac{\text{i}}{2}[S,H_{c}]$ up to second order in $\Lambda$. The effective spin Hamiltonian $\mathcal{H}_{\text{eff}}$ is found by integrating out the $\gamma$–fermions that diagonalize $H_{\text{SC}}$:
@@
The first step is to find a generator $S$ such that $[S,H_{\text{SC}}]=\text{i}H_{c}$ is satisfied. The coupling term is rewritten in terms of the $\gamma$ operators,
@@
The prefactors are determined by comparing the left-hand side of the equation $[S,H_{\text{SC}}]=\text{i}H_{c}$ to the right-hand side. This gives expressions of the kind
@@
$L_{i,j}$ $\frac{\Lambda^{2}}{4N^{2}}\cos\left(\Delta{Q}^{\uparrow\downarrow}\cdot({r}_{i}+{r}_{j})\right)\sum_{{q}}A^{(2)}_{{q},\downarrow,\uparrow}\cos({q}\cdot{\tau})$
$M_{i,j}$ $-\frac{\Lambda^{2}}{4N^{2}}\sin\left(\Delta{Q}^{\uparrow\downarrow}\cdot({r}_{i}+{r}_{j})\right)\sum_{{q}}A^{(2)}_{{q},\downarrow,\uparrow}\cos({q}\cdot{\tau})$ $M_{i,j}=0$ when $\Delta Q^{\uparrow,\downarrow}=0$
@@
${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(5)}$ $-u_{{k}+{q},\sigma}u^{*}_{{k},\sigma^{\prime}}v^{*}_{-{k}-{q},\sigma}v_{-{k},\sigma^{\prime}}\frac{n(\tilde{E}_{{k}+{q},\sigma})-n(\tilde{E}_{{k},\sigma^{\prime}})}{\tilde{E}_{{k}+{q},\sigma}-\tilde{E}_{{k},\sigma^{\prime}}}$ ${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(5)}=({F}_{{k}+q,-{q},\sigma^{\prime},\sigma}^{(5)})^{*}$
${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(6)}$ $-u_{{k}+{q},\sigma}v_{{k},\sigma^{\prime}}v^{*}_{-{k}-{q},\sigma}u^{*}_{-{k},\sigma^{\prime}}\frac{n(\tilde{E}_{{k}+{q},\sigma})+n(\tilde{E}_{-{k},\sigma^{\prime}})-1}{\tilde{E}_{{k}+{q},\sigma}+\tilde{E}_{-{k},\sigma^{\prime}}}$ ${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(6)}=({F}_{{k}+q,-{q},\sigma^{\prime},\sigma}^{(7)})^{*}$
${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(7)}$ $-v^{*}_{{k}+{q},\sigma}u^{*}_{{k},\sigma^{\prime}}u_{-{k}-{q},\sigma}v_{-{k},\sigma^{\prime}}\frac{n(\tilde{E}_{-{k}-{q},\sigma})+n(\tilde{E}_{{k},\sigma^{\prime}})-1}{\tilde{E}_{-{k}-{q},\sigma}+\tilde{E}_{{k},\sigma^{\prime}}}$ ${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(7)}=({F}_{{k}+q,-{q},\sigma^{\prime},\sigma}^{(6)})^{*}$
${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(8)}$ $-v^{*}_{{k}+{q},\sigma}v_{{k},\sigma^{\prime}}u_{-{k}-{q},\sigma}u^{*}_{-{k},\sigma^{\prime}}\frac{n(\tilde{E}_{-{k}-{q},\sigma})-n(\tilde{E}_{-{k},\sigma^{\prime}})}{\tilde{E}_{-{k}-{q},\sigma}-\tilde{E}_{-{k},\sigma^{\prime}}}$ ${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(8)}=({F}_{{k}+q,-{q},\sigma^{\prime},\sigma}^{(8)})^{*}$
@@ I.3 Numerical calculation of the coupling constants
The coupling constants are calculated on a lattice with $N=10000$ lattice points. We have confirmed that this size is sufficient for the coefficients to converge. The exception is when the magnitude of the superconducting OP is small, because then the derivative of the Fermi-Dirac distribution is spiky and a larger number of lattice points is required.
@@
The coupling constants belonging to Fig. 1(c) and 2(c) in the main text are shown in Fig. S1. We observe that the coupling strength generally decreases when the distance $\tau=r_{j}-r_{i}$ increases. $J_{i,j}$ is an order of magnitude greater than the other coupling constants, and wants the spins to be aligned in a collinear configuration. The coupling constants $L_{i,j}$ and $M_{i,j}$ depend on the absolute position $r_{i}$ of the spins. The coupling constants in Fig. S1 are used to calculate the ground state configuration. Similar calculations were performed to calculate the coupling constants for the other parameter configurations for which we calculate the ground state.
@@
Figure S1: The coupling constants in units of $\Lambda^{2}/t$ for a superconductor with equal-spin Cooper pairs. The chemical potential is $\mu=-t$ and the magnitude of the superconducting OP is $\Delta^{\sigma}=0.1t$. A spin current $aQ^{\sigma}=2\pi\sigma\cdot 7/1000$ flows through the superconductor. The coupling constants shown here were used to calculate the ground state configuration in Fig. 2(c) in the main text.
@@ II Calculating the ground state configuration
The ground state configuration is found ”brute force” by calculating $\mathcal{H}_{\text{eff}}$ for $M$ equally spaced values for $\varphi\in[0,2\pi)$ and $M/2$ equally spaced values for $\theta\in[0,\pi)$, and finding the configuration that gives the lowest value for $\mathcal{H}_{\text{eff}}$. This gives in total $M^{4}/4$ combinations for the spins $\bm{S}_{1}$ and $\bm{S}_{2}$. $M=360$ gives a 1 degree uncertainty, and this is the value we have used when determining the ground state.
The ground state configuration remains non-collinear and center-of-mass dependent, as in the main manuscript, when the chemical potential is increased or when the spin current magnitude is decreased. This is shown in Fig. S2.
Upon decreasing the magnitude of the superconducting OP, the coupling constant $J_{i,j}$ becomes dominant and the ground state configuration becomes approximately collinear without specific preferences for the axis alignment.
@@
Figure S2: The ground state configuration in the $xy$-plane of two spins $\bm{S}_{1}$ and $\bm{S}_{2}$ whose interaction is mediated by a superconductor with equal-spin Cooper pairs. The parameters used are the same as in Fig. 2(c) in the main text, except (a) $\mu/t=+1$, (b) $aQ^{\sigma}=2\pi\sigma\cdot 3/1000$ and (c) $aQ^{\sigma}=2\pi\sigma\cdot 5/1000$.
III Magnon gap
III.1 Equal spin pairing

### v2

@@ Abstract
The fact that the intrinsic magnon gap in the antiferromagnetic insulator is enhanced by supercurrent for one magnon species but suppressed for the other lifts their degeneracy. Interestingly, this is the key to inducing magnon spin currents in antiferromagnets via for instance the spin Seebeck effect Rezende et al. (2019). Our results thus demonstrate a route to electrically tunable magnon spin currents in antiferromagnetic insulators, controlled by a dissipationless supercurrent. With regard to the magnitude of the supercurrent-induced magnon gap, our approach is perturbative in nature and thus only small modifications of the magnon gap are within the regime of validity of our methodology. However, the physical mechanism, i.e., the supercurrent-induced spin polarization in the superconductor, is present even in the strong-coupling regime. Thus, a non-perturbative approach using stronger coupling would very likely increase this magnitude and leave the qualitative dependence of the magnon gap on supercurrent similar as that in the weak-coupling regime.
@@ Acknowledgements.
A. Di Bernardo, C. Degen, S. Kölling, K. Knapp, U. Ognjanovic, M. Amundsen, and N. Banerjee are thanked for useful discussions. This work was supported by the Research Council of Norway through Grants No. 353894, 323766 and its Centres of Excellence funding scheme Grant No. 262633 “QuSpin.” The numerical calculations were performed on resources provided by Sigma2—the National Infrastructure for High Performance Computing and Data Storage in Norway, project NN9577K.
@@ I.1 Diagonalizing the Hamiltonian
The mean-field Hamiltonian for a supercurrent-carrying triplet superconductor with equal-spin Cooper pairs is
@@ I.2 Effective spin-spin interaction
The coupling between the itinerant electrons in the superconductor and the classical spins is given by
@@
The goal is to perform a Schrieffer-Wolff transformation to get an effective spin model. The full Hamiltonian is $H=H_{c}+H_{\text{SC}}$, where we treat $H_{c}$ as a perturbation. Note that diagonal elements of the perturbation $H_{c}$ should be absorbed into $H_{\text{SC}}$ in order to allow for first-order contributions in the transformed Hamiltonian. In other words, if $\langle H_{c}\rangle_{\gamma}\neq 0$, these terms should be included in $H_{\text{SC}}$. $\langle H_{c}\rangle_{\gamma}$ is calculated in Eq. (S21), and in the presence of a pure charge or spin supercurrent, this term is zero. For a spin-polarized charge supercurrent, the first-order term, which is effectively a magnetic field, will dominate the spin Hamiltonian. From here on, we assume that the interaction between the magnetic adatoms is induced by a pure charge or spin supercurrent, and we calculate the spin-spin interaction to second order in $\Lambda$. The Hamiltonian is canonically transformed to $H^{\prime}=\text{e}^{\text{i}S}H\text{e}^{-\text{i}S}$, where $S$ is a hermitian operator known as a generator. This transformed Hamiltonian has the same spectrum as the original Hamiltonian. If the generator is chosen such that $S,H_{\text{SC}}]=\text{i}H_{c}$, the transformed Hamiltonian is $H^{\prime}=H_{\text{SC}}+\frac{\text{i}}{2}[S,H_{c}]$ up to second order in $\Lambda$. The effective spin Hamiltonian $\mathcal{H}_{\text{eff}}$ is found by integrating out the $\gamma$–fermions that diagonalize $H_{\text{SC}}$:
@@
The first step is to find a generator $S$ such that $S,H_{\text{SC}}]=\text{i}H_{c}$ is satisfied. The coupling term is rewritten in terms of the $\gamma$ operators,
@@
The prefactors are determined by comparing the left-hand side of the equation $S,H_{\text{SC}}]=\text{i}H_{c}$ to the right-hand side. This gives expressions of the kind
@@
$L_{i,j}$ $\frac{\Lambda^{2}}{4N^{2}}\cos\left(\Delta{Q}^{\uparrow\downarrow}\cdot({r}_{i}+{r}_{j})\right)\sum_{{q}}A^{(2)}_{{q},\downarrow,\uparrow}\cos({q}\cdot{\tau})$ -
$M_{i,j}$ $\frac{\Lambda^{2}}{4N^{2}}\sin\left(\Delta{Q}^{\uparrow\downarrow}\cdot({r}_{i}+{r}_{j})\right)\sum_{{q}}A^{(2)}_{{q},\downarrow,\uparrow}\cos({q}\cdot{\tau})$ $M_{i,j}=0$ when $\Delta Q^{\uparrow,\downarrow}=0$
@@
${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(5)}$ $u_{{k}+{q},\sigma}u^{*}_{{k},\sigma^{\prime}}v^{*}_{-{k}-{q},\sigma}v_{-{k},\sigma^{\prime}}\frac{n(\tilde{E}_{{k}+{q},\sigma})-n(\tilde{E}_{{k},\sigma^{\prime}})}{\tilde{E}_{{k}+{q},\sigma}-\tilde{E}_{{k},\sigma^{\prime}}}$ ${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(5)}=({F}_{{k}+q,-{q},\sigma^{\prime},\sigma}^{(5)})^{*}$
${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(6)}$ $u_{{k}+{q},\sigma}v_{{k},\sigma^{\prime}}v^{*}_{-{k}-{q},\sigma}u^{*}_{-{k},\sigma^{\prime}}\frac{n(\tilde{E}_{{k}+{q},\sigma})+n(\tilde{E}_{-{k},\sigma^{\prime}})-1}{\tilde{E}_{{k}+{q},\sigma}+\tilde{E}_{-{k},\sigma^{\prime}}}$ ${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(6)}=({F}_{{k}+q,-{q},\sigma^{\prime},\sigma}^{(7)})^{*}$
${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(7)}$ $v^{*}_{{k}+{q},\sigma}u^{*}_{{k},\sigma^{\prime}}u_{-{k}-{q},\sigma}v_{-{k},\sigma^{\prime}}\frac{n(\tilde{E}_{-{k}-{q},\sigma})+n(\tilde{E}_{{k},\sigma^{\prime}})-1}{\tilde{E}_{-{k}-{q},\sigma}+\tilde{E}_{{k},\sigma^{\prime}}}$ ${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(7)}=({F}_{{k}+q,-{q},\sigma^{\prime},\sigma}^{(6)})^{*}$
${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(8)}$ $v^{*}_{{k}+{q},\sigma}v_{{k},\sigma^{\prime}}u_{-{k}-{q},\sigma}u^{*}_{-{k},\sigma^{\prime}}\frac{n(\tilde{E}_{-{k}-{q},\sigma})-n(\tilde{E}_{-{k},\sigma^{\prime}})}{\tilde{E}_{-{k}-{q},\sigma}-\tilde{E}_{-{k},\sigma^{\prime}}}$ ${F}_{{k},{q},\sigma,\sigma^{\prime}}^{(8)}=({F}_{{k}+q,-{q},\sigma^{\prime},\sigma}^{(8)})^{*}$
@@ I.3 Numerical calculation of the coupling constants
We calculate the coupling constants numerically at zero temperature. The coupling constants can also be calculated at higher temperatures, but the spin ground state of the adatoms is only defined at $T=0$. However, in the numerical procedure used to calculate the coupling constants in $\mathcal{H}_{\text{eff}}$, it is convenient to include a small temperature $T=0.01T_{c}$ well below the superconducting critical temperature $T_{c}$ to effectively turn the Fermi-Dirac distribution functions into slightly smeared step functions. This facilitates numerical convergence and avoids the requirement of computationally infeasible system sizes. The coupling constants are calculated on a lattice with N=10  [^0] N=10\;000 lattice points. We have confirmed that this size is sufficient for the coefficients to converge as a function of $N$. The exception is when the magnitude of the superconducting OP is small, because then the derivative of the Fermi-Dirac distribution is spiky and a larger number of lattice points is required.
@@
The coupling constants belonging to Fig. 1(c) and 2(c) in the main text are shown in Fig. S1. We observe that the coupling strength generally decreases when the distance $\tau=r_{j}-r_{i}$ increases. The coupling constants $L_{i,j}$ and $M_{i,j}$ depend on the absolute position $r_{i}$ of the spins. The numerically calculated coupling constants that were used to determine the ground state configuration for the other parameter variations (no supercurrent, charge supercurrent, different supercurrent strengths) are not shown here.
@@
Figure S1: The coupling constants in units of $\Lambda^{2}/t$ for a superconductor with equal-spin Cooper pairs. The chemical potential is $\mu=-t$ and the magnitude of the superconducting OP is $\Delta^{\sigma}=0.1t$. A spin current $Q^{\sigma}/Q_{c}\approx 0.87\sigma$, or equivalently $aQ^{\sigma}=2\pi\sigma\cdot 7/1000$, flows through the superconductor. The coupling constants shown here were used to calculate the ground state configuration in Fig. 2(c) in the main text.
I.4 Higher dimensions
Our model is straightforwardly generalized to 2D or 3D by summing over all the components of the vectors $\bm{k}$ and $\bm{q}$ in the momentum space sums in Table 1. Additionally, the expressions for $\epsilon_{\bm{k}}$ and $\Delta_{\bm{k}}^{\sigma}$ in any dimension are $\epsilon_{\bm{k}}=-\mu-t\sum_{\bm{\delta}}e^{i\bm{k}\cdot\bm{\delta}}$ and $\Delta_{\bm{k}}^{\sigma}=\sum_{\delta}\Delta_{\delta}^{\sigma}e^{i\bm{k}\cdot\bm{\delta}}$, where $\bm{\delta}$ refers to the vectors to all nearest neighbors. The main obstacle is that numerical calculations in 2D and 3D become extremely time-consuming since all the coupling constants contain terms of the kind $(n(E_{1})-n(E_{2}))/(E_{1}-E_{2})$. At low temperatures, this resembles a Dirac delta function for $\bm{k}$-values where $E_{1}\rightarrow E_{2}$, and a large system size is therefore needed to sample this part. For this reason, we have focused on the 1D case, but do not expect any qualitative changes to the effective spin-spin Hamiltonian in higher dimensions. The fact that our main results hold both in 2D and in 3D can be proven analytically as follows. Consider for instance the Dzyaloshinskii-Moriya interaction. We have
$D_{i,j}=\frac{\Lambda^{2}}{4N^{2}}\sum_{\bm{q}}A^{(1)}_{\bm{q}+\Delta\bm{Q}^{\uparrow\downarrow},\downarrow,\uparrow}\sin(\bm{q}\cdot\bm{\tau}),$ (S14)
where $\sum_{\bm{q}}$ now runs over a 2D or 3D Brillouin zone, $\Delta\bm{Q}^{\uparrow\downarrow}=\bm{Q}^{\uparrow}-\bm{Q}^{\downarrow}$, and in general $A^{(1)}_{\bm{q},\downarrow,\uparrow}=A^{(1)}_{-\bm{q},\uparrow,\downarrow}$. When the magnitude of the superconducting order parameter is spin independent, $A^{(1)}_{\bm{q},\downarrow,\uparrow}=A^{(1)}_{\bm{q},\uparrow,\downarrow}$, so $A_{\bm{q}}^{(1)}$ is symmetric in $\bm{q}$. When $\Delta\bm{Q}^{\uparrow\downarrow}=0$, the DM coefficient is a sum over an even function times an odd function, which gives $D_{i,j}=0$. When $\Delta\bm{Q}^{\uparrow\downarrow}\neq 0$, we have $A^{(1)}_{\bm{q}+\Delta\bm{Q^{\uparrow\downarrow}},\downarrow,\uparrow}\neq A^{(1)}_{-\bm{q}+\Delta\bm{Q}^{\uparrow\downarrow},\downarrow,\uparrow}$, so the DM term is nonzero. The larger the spin current is, the further $A^{(1)}_{\bm{q}+\Delta\bm{Q}^{\uparrow\downarrow},\downarrow,\uparrow}$ is shifted from being even and the larger $D_{i,j}$ becomes. The center-of-mass dependent coefficients $L_{i,j}$ and $M_{i,j}$ are given by
$L_{i,j}=\frac{\Lambda^{2}}{4N^{2}}\cos\left(\Delta\bm{Q}^{\uparrow\downarrow}\cdot(\bm{r}_{i}+\bm{r}_{j})\right)\sum_{\bm{q}}A^{(2)}_{\bm{q},\downarrow,\uparrow}\cos(\bm{q}\cdot\bm{\tau})$ (S15)
$M_{i,j}=-\frac{\Lambda^{2}}{4N^{2}}\sin\left(\Delta\bm{Q}^{\uparrow\downarrow}\cdot(\bm{r}_{i}+\bm{r}_{j})\right)\sum_{\bm{q}}A^{(2)}_{\bm{q},\downarrow,\uparrow}\cos(\bm{q}\cdot\bm{\tau}).$ (S16)
Since the magnitude of the superconducting order parameter is taken to be spin independent and real in $\bm{k}$-space, we have $A^{(2)}_{\bm{q},\downarrow,\uparrow}=A^{(2)}_{-\bm{q},\downarrow,\uparrow}$. The sum is therefore a sum over two even functions in $\bm{q}$, which is generally nonzero. In fact, the $\bm{q}$-sum is identical to the $\bm{q}$-sum in the anisotropic exchange coupling $K_{i,j}$ when the magnitude of the superconducting order parameter is spin independent and real in $\bm{k}$-space. In the absence of a spin supercurrent, $M_{i,j}=0$ while $L_{i,j}\neq 0$, although $L_{i,j}$ loses its center-of-mass dependence. In the presence of a spin supercurrent, both coupling constants are generally nonzero and dependent on the center-of-mass coordinate regardless of the dimensionality of the system.
@@ II Calculating the ground state configuration
The ground state configuration is found ”brute force” by calculating $\mathcal{H}_{\text{eff}}$ for $M$ equally spaced values for $\varphi\in[0,2\pi)$ and $M/2$ equally spaced values for $\theta\in[0,\pi)$, and finding the configuration that gives the lowest value for $\mathcal{H}_{\text{eff}}$. This gives in total $M^{4}/4$ combinations for the spins $\bm{S}_{1}$ and $\bm{S}_{2}$. $M=360$ gives a 1 degree uncertainty, and this is the value we have used when determining the ground state of the dimers (two spins). When we calculate the ground state of trimers (three spins), we use $M=60$ which gives an uncertainty of 6 degrees. The reduced accuracy for the trimers is due to the increased computational cost of calculating the ground state for three spins.
The ground state configuration remains non-collinear and center-of-mass dependent, as in the main manuscript, when the chemical potential is increased or when the spin current magnitude is decreased. This is shown in Fig. S2. In Fig. S3, the ground state dependence on the spin and charge supercurrent magnitude is shown. We see that the ground state configuration can be tuned by a supercurrent. Upon decreasing the magnitude of the superconducting OP, the coupling constant $J_{i,j}$ becomes dominant and the ground state configuration becomes approximately collinear without specific preferences for the axis alignment. Since the effective Hamiltonian contains five different types of interactions which are all generally non-zero in the presence of a spin supercurrent, there are competing interactions in the Hamiltonian which cannot be minimized simultaneously. This is exactly the definition of magnetic frustration. Therefore, we expect that the magnetic ground state is in general frustrated and that the spin supercurrent does tune the frustration by altering the relative magnitude of the coefficients.
@@
Figure S2: The ground state configuration in the $xy$-plane of two spins $\bm{S}_{1}$ and $\bm{S}_{2}$ whose interaction is mediated by a superconductor with equal-spin Cooper pairs. The parameters used are the same as in Fig. 2(c) in the main text, except (a) $\mu/t=+1$, (b) $Q^{\sigma}/Q_{c}\approx 0.37\sigma$ and (c) $Q^{\sigma}/Q_{c}\approx 0.62\sigma$.
Figure S3: (a-b): Spin supercurrent dependence on the magnetic ground state. The position of $\bm{S}_{1}$ is $r_{1}/a=0$ in (a) and $r_{1}/a=8$ in (b). (c) Charge supercurrent dependence on the magnetic ground state. The ground state at $Q=0$ is degenerate at $\tau/a=3$ and 6, but the energy is not minimized for the ferromagnetic alignment along $y$.
Figure S4 shows the magnetic ground state for spin trimers with different spin positions and supercurrents in the superconductor. Importantly, the results from the dimer calculation hold even for the trimers. First, we see that we get non-collinear ground states. In the absence of a spin supercurrent, this non-collinearity arises due to antiferromagnetic next-nearest neighbor coupling with almost the same strength as the nearest neighbor antiferromagnetic coupling. When the spin supercurrent is turned on, we get DM interaction between the spins as an additional source of spin non-collinearity. Second, the ground state of the trimer is center-of-mass dependent in the presence of a spin supercurrent. Therefore, the magnetic ground state of the impurity spins can be tuned by placing the spins appropriately on the surface. Moreover, the ground state of the trimer can be tuned with the magnitude of the spin supercurrent. This is shown in Fig. 2(e) in the main text.
Figure S4: Ground states for spin trimers with the three spins placed at various positions on top of the superconductor. The spins are in the xyxy-plane, and the position of each of the three spins is shown along the horizontal axis. The magnitude of the charge supercurrent is given by $Q^{\sigma}/Q_{c}\approx 0.87$, and the magnitude of the spin supercurrent is given by $Q^{\sigma}/Q_{c}\approx 0.87\sigma$ as in the main paper. The accuracy of the spin orientations is 6 degrees (limited by a high numerical computation cost).
III The origin for the center-of-mass coordinate of the magnetic adatoms
The physical quantity that defines the origin of the axis from which the position $r_{i}$ is measured is the phase difference between the superconducting order parameters describing Cooper pairs with opposite spin. The superconducting order parameter for spin-$\sigma$ Cooper pairs is
$\Delta_{i,j}^{\sigma}=\Delta_{\delta}^{\sigma}\text{e}^{\text{i}Q^{\sigma}\cdot({r}_{i}+{r}_{j})},$ (S18)
where $i,j$ are nearest neighbors. Now, consider $r_{i}<r_{j}$ without loss of generality. The phase difference between spin up and spin down Cooper pairs is then $\Delta Q(2r_{i}+a)$. The origin $r=0$ is the point where the phase difference between the Cooper pairs is $\Delta Qa$ when the electrons in the Cooper pairs are located at this lattice site and the one to the right of it. For $r_{j}<r_{i}$, the order parameter picks up a phase $\pi$ due to $\Delta_{\delta}^{\sigma}$ switching sign, but this does not affect the phase difference between the order parameters describing Cooper pairs with opposite spin. Since the phase-winding causes the supercurrent-carrying superconducting order parameters $\Delta^{\sigma}_{i,j}$ to be periodic in space, there are several equivalent positions that may be used to define the point $r_{i}=0$. Measuring this phase difference directly would be difficult, but it nevertheless has an experimentally measurable consequence: the center-of-mass dependence of the spin interactions predicted in our work as seen by experimentally by placing several pairs of spins at an arbitrary position on the surface of a superconductor. The magnetic ground-state of the spin pairs will then in general be different precisely due to the center-of-mass dependence induced by the spin supercurrent in our system.
IV Magnon gap
IV.1 Equal spin pairing